OpenStudy (anonymous):

Solve for a: w=(2πa^2)+(2πah)

OpenStudy (anonymous):

I can't find a way to solve this. Are you sure this is solvable?

OpenStudy (anonymous):

Not sure, my teacher assigned it for review but i cannot find an answer. Maybe he made a typo

OpenStudy (anonymous):

\[w\text{=}\left(2\pi a^2\right)+(2\pi a h) \]\[\left\{a=\frac{-h \pi -\sqrt{\pi } \sqrt{h^2 \pi +2 w}}{2 \pi },a=\frac{-h \pi +\sqrt{\pi } \sqrt{h^2 \pi +2 w}}{2 \pi }\right\} \]